A scientist determined that the bones from a mastodon had lost 70.2% of their carbon-14.How old were the bones at the time they were​ discovered?
Found 2 solutions by ikleyn, Boreal:
Answer by ikleyn(52787)

.
A scientist determined that the bones from a mastodon had lost 70.2% of their carbon-14.
How old were the bones at the time they were discovered?
~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~~

See the lesson Using logarithms to solve real world problems in this site.

You will find there the description of the carbon dating method and the sample for your problem to solve.


Answer by Boreal(15235)

It is [ln(.298)/-0.693]*5730=1.747*5730=10010.2 years
-0.693 is ln (1/2), the half life, or 50%. If 75% were lost, that would be 2 half-lives or 11460 years, so this is in the range.
Determine the number of half life (1/2)^n=0.298
ln of both sides is n ln(1/2)=ln 0.298. Divide both sides by ln (1/2), which is -0.693